OpenStudy (anonymous):

the formula r=2√5L can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet how far will a car skid at 70 mph?

OpenStudy (anonymous):

245

OpenStudy (anonymous):

L=5*(r/2)^2

jimthompson5910 (jim_thompson5910):

\[\large r=2\sqrt{5L}\] \[\large 70=2\sqrt{5L}\] \[\large \frac{70}{2}=\sqrt{5L}\] \[\large 35=\sqrt{5L}\] \[\large 35^2=5L\] \[\large 1225=5L\] \[\large \frac{1225}{5}=L\] \[\large 245=L\] \[\large L=245\] So the car will skip 245 ft.

jimthompson5910 (jim_thompson5910):

lol oops meant to say skid instead of skip, that's one hopped up car